# A toric degeneration of Kronecker moduli spaces

### v1

@@ 1. Introduction
$\theta:=(-r_{2},r_{1})\in\mathbb{Z}^{2}\cong\chi(G).$
The Kronecker moduli space $\mathrm{K}^{n}_{r_{1}r_{2}}$ is the projective geometric invariant theory (GIT) quotient
@@
The basic steps in constructing the Gelfand–Cetlin toric degeneration of the Grassmannian $\operatorname{\mathrm{Gr}}(r,n)$ are as follows. One first fixes a term order on the polynomial ring $\mathbb{C}[x_{ij}]$, where the $x_{ij}$ are the coordinates on an $r\times n$ matrix, such that the leading monomial $\mathrm{LM}(m)$ of any minor $m$ is the diagonal monomial. The coordinate ring $\mathbb{C}[\operatorname{\mathrm{Gr}}(r,n)]$ of the Grassmannian is a sub-algebra of $\mathbb{C}[x_{ij}]$. Then:
@@
    $\{\text{SSYT of size }r\times k:k\in\mathbb{N}\},$
@@
    Describe a finite set $S$ of SSYT such that the associated subset of $\mathcal{B}$ is a SAGBI basis for $\mathbb{C}[\operatorname{\mathrm{Gr}}(r,n)]$; that is, for any $f\in\mathbb{C}[\operatorname{\mathrm{Gr}}(r,n)]$, the leading term of $f$ is a monomial in the leading terms of elements from this subset.
@@
To state our results more precisely, we recall the definition of linked tableau pairs. Fix an $a\in\frac{1}{\gcd(r_{1},r_{2})}\mathbb{N}.$ Let $\alpha^{-}$ be the partition of shape $r_{1}\times ar_{2}$ and $\alpha^{+}$ the partition of shape $r_{2}\times ar_{1}$. Let $T^{-}$ be a tableau of shape $\alpha^{-}$, filled with entries of the form $ij$, where
@@
The pair $(T^{+},T^{-})$ form a linked tableaux pair if there is a bijection between the labels of $T^{-}$ and the labels of $T^{+}$ taking a label $iq$ in the $p^{th}$ row of $T^{-}$ to a label $ip$ in the $q^{th}$ row of $T^{+}.$ We call this bijection the link between the pair. Define an ordering on the labels by setting $ip<jq$ if $i<j$ or $i=j$ and $p<q.$ In this way, we extend the notion of semi-standard to these tableaux with double entries: a linked pair is semi-standard if both $T^{+}$ and $T^{-}$ are semi-standard. A semi-standard linked pair is primitive, if it cannot be subdivided into two sets of semi-standard linked pairs (see Definition 2.8 for a precise definition).
@@
This means that the set of primitive linked tableaux index a SAGBI basis. As an application, we obtain the following statement:
@@ 2.1. SAGBI bases and toric degenerations
Let $A\subset\mathbb{C}[x_{1},\dots,x_{n}]$ be a sub-algebra. Given a grading $\mathrm{gr}$, let $\mathrm{in}_{<}(f)$ be the initial term (or terms) of $f\in A$. Then a set $f_{1},\dots,f_{k}$ is a SAGBI basis if for each $f_{i}$, $\mathrm{in}_{<}(f_{i})$ is a monomial and
@@
One reason to be interested in SAGBI bases is their role in toric degenerations. Let $X$ be a projective variety $\operatorname{\mathrm{Proj}}(R)$, for $R$ a graded sub-algebra of a sufficiently large polynomial ring $\mathbb{C}[x_{1},\dots,x_{N}]$. Fix a term order $\mathrm{in}_{<}$ on $\mathbb{C}[x_{1},\dots,x_{N}]$, and suppose that $f_{1},\dots,f_{k}$ is a finite SAGBI basis for $R$. Then $\mathrm{in}_{<}(R)$ is generated by the $\mathrm{in}_{<}(f_{i})$ (in particular, it is finitely generated). There is a flat family of algebras $\mathcal{R}\to\mathbb{C}$ such that $\mathcal{R}_{1}=R$ and $\mathcal{R}_{0}=\mathrm{in}_{<}(R)$, obtained by deforming the polynomials in $R$ to their initial terms [10].
The initial algebra $\mathrm{in}_{<}(R)$ is the coordinate ring of the (possibly) non-normal toric variety $\operatorname{\mathrm{Spec}}(\mathrm{in}_{<}(R))$. We will be most interested in the projective version of this deformation, and in the case where the monoid $\mathrm{in}_{<}(R)$ is the monoid associated to a rational cone. Let $\mathrm{exp}(m)$ of a monomial $m\in\mathbb{C}[x_{1},\dots,x_{N}]$ be the exponent vector. We take the cone $C$ over the elements
($(\{(\mathrm{exp}(\mathrm{in}_{<}(f_{i})),\mathrm{gr}(f_{i}))\in\mathbb{R}^{N}\times\mathbb{R}:i=1,\dots,k\}.$
@@
$C\cap L=\{(\mathrm{exp}(m),\mathrm{gr}(m)):m\text{ a monomial in }\mathrm{in}_{<}(R)\}.$
@@
This appears in various places in the literature; some relevant (but not exhaustive) citations include [10, Theorem 15.17] (flatness, including in the case where $\mathrm{in}_{<}$ is not a total order) and [26]. In [19] there is a nice description of the projective version.
We now identify the projective toric variety $\operatorname{\mathrm{Proj}}(\mathrm{in}_{<}(R))$. Let $C$ be as above, let $P$ be the polytope obtained by taking the height one slice of $C$. Note that $P$ may not be a lattice polytope, and also is not necessarily full dimensional. By considering $P$ in the subspace spanned by itself, we can define a normal fan $F$.
@@
Proof.
We first assume that $\mathrm{in}_{<}(R)$ is generated in degree $1$, so that $C$ is generated by $C\cap\mathbb{Z}^{N}\times\{1\}.$ Then the monoid satisfies the assumptions of Lemma 2.3 in [1], so $\operatorname{\mathrm{Proj}}(\mathrm{in}_{<}(R))=\operatorname{\mathrm{Proj}}(\mathbb{C}[C])$ is the normal toric variety associated to the polytope $P^{\prime}$, defined so that $P^{\prime}\times\{1\}=1$; i.e. $X_{F}$ as required.
If $\mathrm{in}_{<}(R)$ is not generated in degree $1$, there is some multiple $mP$ of $P$ such that $mP$ is a lattice polytope. We can consider the monoid $\oplus_{j}(\mathrm{in}_{<}(R))_{jm}$, and divide our original grading by $m$. $\operatorname{\mathrm{Proj}}$ of this monoid is isomorphic to the original; we have just taken the $m^{th}$ Veronese; i.e. changed the polarization. Then the new cone $\tilde{C}$ satisfies the requirement that $\tilde{C}$ is generated by $\tilde{C}\cap\mathbb{Z}^{N}\times\{1\}$, so again, we can conclude that $\operatorname{\mathrm{Proj}}(\oplus_{j}(\mathrm{in}_{<}(R))_{jm})$ is the normal toric variety associated to the polytope given by the height one slice of $\tilde{C}$. But this polytope is isomorphic to $mP$, and thus has the same normal fan. ∎
That is, the affine cone over the toric variety $X_{F}$ is the normal affine toric variety $\operatorname{\mathrm{Spec}}(\mathrm{in}_{<}(R))$.
@@ 2.2. Construction of Kronecker moduli
We denote the quiver with $n$ arrows as $Q^{n}$, and set $Q^{n}_{0}=\{1,2\}$ to be the set of vertices, and $Q^{n}_{1}$ the set of arrows.
@@
$(x^{i}_{jk})_{\begin{subarray}{c}1\leq j\leq r_{2},\\ 1\leq k\leq r_{2}\end{subarray}}$
@@
The group $G:=\operatorname{\mathrm{GL}}(r_{1})\times\operatorname{\mathrm{GL}}(r_{2})$ acts by change of basis on $\operatorname{\mathrm{Rep}}(Q^{n},\mathbf{r})$. Let $\lambda_{0}:=\mathbb{C}^{*}\to G$ be the diagonal one parameter subgroup. Then the image of $\lambda_{0}$ acts trivially on $G$. For any character $\theta\in\chi(G)$ such that $\langle\theta,\lambda_{0}\rangle=0$, we can define the GIT quotient
@@ 2.3. Semi-invariants of Kronecker moduli
Fix a Kronecker moduli space, $\mathrm{K}^{n}_{r_{1}r_{2}}$, and a character $\alpha\in\chi(G)$ satisfying $\langle\alpha,\lambda_{0}\rangle=0.$ That is, if $\alpha=(-\alpha_{1},\alpha_{2})$, using the identification $\chi(G)\cong\mathbb{Z}^{2}$, then
@@
$\mathrm{Mon}(T^{+}):=\prod_{\text{label }ij\text{ in $k^{th}$ row}}x^{i}_{kj},$
$\mathrm{Mon}(T^{-}):=\prod_{\text{label }ik\text{ in $j^{th}$ row}}x^{i}_{kj}.$
@@
Let $\mathrm{LM}(f)$ denote the leading monomial of an element of the semi-invariant ring, under any term order such that the leading monomial of any maximal minor of either
@@
is the diagonal term.
@@ 3. The main theorem
[…]
@@
To complete the proof, we show the following claim, and its analogue for $T^{-}$: $T^{+}$ is semi-standard if, for every two consecutive rows in $T^{+}$, say row $j$ and $j+1$, and every choice of $i,k$,
@@
The claim holds for the following reason. Consider the last label $ik$ in the $j+1$ row. This is the $\sum_{sk\leq ik}v^{s}_{jl}$ entry in this row. Then if $T^{+}$ is semi-standard, the label above $ik$ is at most $ik-1$. This means that the number of labels in row $j$ of form $sk\leq ik-1$ is at least $\sum_{sk\leq ik}v^{s}_{jl}$; that is,
@@
For non-trivial examples, see §5. The simplest example, however, of $C^{n}_{r_{1}r_{2}}$, is when either $r_{1}=1$ or $r_{2}=1$. The cases are isomorphic, we so can consider the cone $C^{n}_{1r}$, and its slice at height one, $P_{r,n}:=P^{n}_{1r}(1)$. In this case, $C^{n}_{1r}$ is generated at height one, and $P_{r,n}$ is the Gelfand–Cetlin polytope (or isomorphic to it, depending on your description). More precisely, the Gelfand–Cetlin polytope is the image of the Gelfand–Cetlin integrable system on the Grassmannian $\operatorname{\mathrm{Gr}}(r,n)$. The Gelfand–Cetlin integrable system can be described directly, or via the SAGBI basis degeneration to the degenerate toric variety $X_{r,n}$ given by SSYT tableaux, as in the motivating example of [14]. That is, in [14], they show that there the integrable system can be obtained by composing the map they produce $\operatorname{\mathrm{Gr}}(r,n)\to X_{r,n}$ with the integrable system on $X_{r,n}$ using its toric structure. As, by definition, $X_{r,n}$ is $\operatorname{\mathrm{Proj}}$ of the monoid obtained by $C^{n}_{1r}$, the image of its integrable system is naturally seen as $P_{r,n}$ as desired.
@@
[…]
@@
In the next theorem, we show that $\mathrm{K}^{n}_{r_{1}r_{2}}$ has a toric degeneration to the toric variety associated to the intersection of (scaled) Gelfand–Cetlin polytopes.
@@
We first show that the intersection in the statement of the theorem makes sense. The lattice polytope $P_{r_{1},nr_{2}}$ lives in the toric lattice $M_{1}:=\operatorname{\mathrm{Mat}}(r_{1},nr_{2};\mathbb{Z})$, and the lattice polytope $P_{r_{2},nr_{1}}$ lives in the lattice $M_{2}:=\operatorname{\mathrm{Mat}}(r_{2},nr_{1};\mathbb{Z})$. We identify both of these lattices with $M=\operatorname{\mathrm{Mat}}(r_{2},r_{1};\mathbb{Z})^{\times n}$, by taking a tuple of matrices in $M$ and either stacking them horizontally (giving an element of $M_{2}$) or vertically and then transposing (giving an element of $M_{1}$).
We then claim that

### v2

@@ 1. Introduction
$\theta:=(-r_{2},r_{1})\in\mathbb{Z}^{2}\cong\chi(G),$
where $\chi(G)$ is the character lattice of $G$. The Kronecker moduli space $\mathrm{K}^{n}_{r_{1}r_{2}}$ is the projective geometric invariant theory (GIT) quotient
@@
The basic steps in constructing the Gelfand–Cetlin toric degeneration of the Grassmannian $\operatorname{\mathrm{Gr}}(r,n)$ are as follows. One first fixes a term order on the polynomial ring $\mathbb{C}[x_{ij}]$, where the $x_{ij}$ are the coordinates on an $r\times n$ matrix, such that the leading monomial $\mathrm{LM}(m)$ of any minor $m$ is the diagonal monomial. The coordinate ring $\mathbb{C}[\operatorname{\mathrm{Gr}}(r,n)]$ of the Grassmannian is a sub-algebra of $\mathbb{C}[x_{ij}]$. Let $\mathrm{SSYT}(n)$ denote the set of semi-standard Young tableaux with labels in the set $\{1,\dots,n\}$. Then:
@@
    $\{\text{rectangular elements of $\mathrm{SSYT}(n)$ with $r$ rows}\},$
@@
    Describe a finite set $S$ in $\mathrm{SSYT}(n)$ such that the associated subset of $\mathcal{B}$ is a SAGBI basis for $\mathbb{C}[\operatorname{\mathrm{Gr}}(r,n)]$; that is, for any $f\in\mathbb{C}[\operatorname{\mathrm{Gr}}(r,n)]$, the leading term of $f$ is a monomial in the leading terms of elements from this subset.
@@
To state our results more precisely, we recall the definition of linked tableau pairs for $\mathrm{K}^{n}_{r_{1}r_{2}}$. We say a partition has shape $p\times q$ if the associated Ferrers’ diagram is the $p\times q$ rectangle. Let $\alpha^{-}$ be the partition of shape $r_{1}\times ar_{2}$, and $\alpha^{+}$ the partition of shape $r_{2}\times ar_{1}$ for some $a\in\frac{1}{\gcd(r_{1},r_{2})}\mathbb{N}.$ Let $T^{-}$ be a tableau of shape $\alpha^{-}$, filled with entries of the form $ij$, where
@@
The pair $(T^{+},T^{-})$ form a linked tableaux pair if there is a bijection between the labels of $T^{-}$ and the labels of $T^{+}$ taking a label $iq$ in the $p^{th}$ row of $T^{-}$ to a label $ip$ in the $q^{th}$ row of $T^{+}.$ We call this bijection the link between the pair. Define an ordering on the labels by setting $ip<jq$ if $i<j$ or $i=j$ and $p<q.$ In this way, we extend the notion of semi-standard to these tableaux with double entries: a linked pair is semi-standard if both $T^{+}$ and $T^{-}$ are semi-standard. A semi-standard linked pair is primitive, if it cannot be subdivided into two sets of semi-standard linked pairs (see Definition 2.8 for a precise definition). For example, the following is a primitive semi-standard linked pair for $r_{1}=2,$ $r_{2}=3$:
[…]
@@
This means that the set of primitive linked tableaux index a SAGBI basis. In this paper, all SAGBI bases are assumed to be finite. As an application, we obtain the following statement:
@@ 2.1. SAGBI bases and toric degenerations
On polynomial rings $\mathbb{C}[x_{1},\dots,x_{n}]$ we consider orders given by weights $w\in\mathbb{R}^{n}_{\geq 0}$. Any such $w$ assigns a weight $w_{i}$ to each $x_{i}$, and a monomial $\prod_{i=1}^{n}x_{i}^{a_{i}}$ is given the weight $\sum_{i=1}^{n}a_{i}w_{i}.$ This allows us to define a weak order on the monomials in the polynomial ring given by their weights. Given a polynomial $f$, $\mathrm{in}_{<}(f)$ is the polynomial made up of the terms of $f$ that achieve the maximum weight among the monomials of $f$. In most cases that we consider, this will be a monomial, which we call the initial term.
@@
Let $A\subset\mathbb{C}[x_{1},\dots,x_{n}]$ be a sub-algebra, together with a choice of weight $w$. Then a set $f_{1},\dots,f_{k}$ is a SAGBI basis if for each $f_{i}$, $\mathrm{in}_{<}(f_{i})$ is a monomial and
@@
One reason to be interested in SAGBI bases is their role in toric degenerations. Let $X$ be a projective variety $\operatorname{\mathrm{Proj}}(R)$, for $R$ a graded sub-algebra of a sufficiently large polynomial ring $\mathbb{C}[x_{1},\dots,x_{N}]$. Suppose that $f_{1},\dots,f_{k}$ is a finite SAGBI basis for $R$ for some choice of weight. Then $\mathrm{in}_{<}(R)$ is generated by the $\mathrm{in}_{<}(f_{i})$ (in particular, it is finitely generated). There is a flat family of algebras $\mathcal{R}\to\mathbb{C}$ such that $\mathcal{R}_{1}=R$ and $\mathcal{R}_{0}=\mathrm{in}_{<}(R)$, obtained by deforming the polynomials in $R$ to their initial terms [9].
The initial algebra $\mathrm{in}_{<}(R)$ is the coordinate ring of the (possibly) non-normal toric variety $\operatorname{\mathrm{Proj}}(\mathrm{in}_{<}(R))$. We will be most interested in the case where the semi-group algebra $\mathrm{in}_{<}(R)$ is the semi-group algebra associated to a rational cone. Let $\mathrm{exp}(m)$ of a monomial $m\in\mathbb{C}[x_{1},\dots,x_{N}]$ be the exponent vector. We take the cone $C$ over the elements
$\{(\mathrm{exp}(\mathrm{in}_{<}(f_{i})),\mathrm{gr}(f_{i}))\in\mathbb{R}^{N}\times\mathbb{R}:i=1,\dots,k\}.$
@@
(1) $C\cap L=\{(\mathrm{exp}(m),\mathrm{gr}(m)):m\text{ a monomial in }\mathrm{in}_{<}(R)\}.$
@@
This appears in various places in the literature; some relevant (but not exhaustive) citations include [9, Theorem 15.17] (flatness, including in the case where $\mathrm{in}_{<}$ is not a total order) and [26]. In [19] there is a nice description of the projective version.
We now identify the projective toric variety $\operatorname{\mathrm{Proj}}(\mathrm{in}_{<}(R))$. We can choose a large enough $m$ such that $P=C\cap\mathbb{R}^{N}\times\{m\}$ is a normal lattice polytope. Let $F$ be the normal fan of $P$; if $P$ is not full-dimensional, we take the normal fan of PxP-x in the real vector subspace given by the span of PxP-x for any lattice point $x$ in $P$. Then we have the following identification:
@@
If we had not assumed (1), then $X_{F}$ would be the normalization of $\operatorname{\mathrm{Proj}}(\mathrm{in}_{<}(R)).$ This is a standard result in toric geometry, but is specifically addressed in the discussion following Proposition 3.4 in [10] in the situation where the cone is not necessarily full dimensional.
@@ 2.2. Construction of Kronecker moduli
We denote the quiver with $n$ arrows as $Q^{n}$, and set $Q^{n}_{0}=\{1,2\}$ to be the set of vertices, and $Q^{n}_{1}$ the set of arrows. There are maps $s,t:Q^{n}_{1}\to Q^{n}_{0}$ taking arrows to their source and target vertices respectively.
@@
(2) $(x^{i}_{jk})_{\begin{subarray}{c}1\leq j\leq r_{2},\\ 1\leq k\leq r_{2}\end{subarray}}$
@@
The group $G:=\operatorname{\mathrm{GL}}(r_{1})\times\operatorname{\mathrm{GL}}(r_{2})$ acts by change of basis on $\operatorname{\mathrm{Rep}}(Q^{n},\mathbf{r})$. The character lattice of $G$, denoted $\chi(G)$, can be identified with $\mathbb{Z}^{2}$, where the standard generators are given by the two determinant maps. Let $\lambda_{0}:=\mathbb{C}^{*}\to G$ be the diagonal one parameter subgroup. Then the image of $\lambda_{0}$ acts trivially on $G$. For any character $\theta\in\chi(G)$ such that $\langle\theta,\lambda_{0}\rangle=0$, we can define the GIT quotient
@@ 2.3. Semi-invariants of Kronecker moduli
Fix a Kronecker moduli space, $\mathrm{K}^{n}_{r_{1}r_{2}}$, and a character $\alpha\in\chi(G)$ satisfying $\langle\alpha,\lambda_{0}\rangle=0.$ That is, if $\alpha=(-\alpha_{1},\alpha_{2})$, then
@@
$\mathrm{Mon}(T^{+}):=\prod_{k=1}^{r_{2}}\prod_{\begin{subarray}{c}\text{labels }ij\\ \text{ in row $k$}\end{subarray}}x^{i}_{kj},$
$\mathrm{Mon}(T^{-}):=\prod_{j=1}^{r_{1}}\prod_{\begin{subarray}{c}\text{labels }ik\\ \text{ in row $j$}\end{subarray}}x^{i}_{kj}.$
@@
Recall that the $A_{i}:=[x^{i}_{jk}]$ are the coordinate matrices defined in (2). Unless otherwise stated, we will let $\mathrm{LM}(f)$ denote the leading monomial of an element of the semi-invariant ring, under any term order that is diagonal. That is, we require that the leading monomial of any maximal minor of either
@@
is the diagonal term. One such term order is given by taking $x^{i}_{jk}>x^{l}_{pq}$ when $i<l$, and within each $A_{i}$, taking the order defined by going across rows left to right and top to bottom.
@@ 3. The main theorem
$C^{n}_{r_{1}r_{2}}:=\mathrm{Cone}(v_{T^{\pm}}:T^{\pm}\text{ is a semi-standard pair}).$
@@
To complete the proof, we show the following claim, and its analogue for $^{-}: T+T^{+}$ is semi-standard if, for every two consecutive rows in $T^{+}$, say row $j$ and $j+1$, and every choice of $i,k$,
@@
The claim holds for the following reason. Consider the last label $ik$ in the $j+1$ row. This is the $\sum_{sk\leq ik}v^{s}_{jl}$ entry in this row. Then if $T^{+}$ is semi-standard, the label above $ik$ is at most ik1ik-1. This means that the number of labels in row $j$ of form $sk\leq ik-1$ is at least $\sum_{sk\leq ik}v^{s}_{jl}$; that is,
@@
For non-trivial examples, see §5. The simplest example, however, of $C^{n}_{r_{1}r_{2}}$, is when either $r_{1}=1$ or $r_{2}=1$. The cases are isomorphic, we so can consider the cone $C^{n}_{1r}$, and its slice at height one, $P_{r,n}:=P^{n}_{1r}(1)$. In this case, $C^{n}_{1r}$ is generated at height one, and $P_{r,n}$ is the Gelfand–Cetlin polytope (or isomorphic to it, depending on your description). In [14], they show that there the integrable system can be obtained by composing the map they produce from $\operatorname{\mathrm{Gr}}(r,n)\to X_{r,n}$ with the integrable system on $X_{r,n}$ using its toric structure. As, by definition, $X_{r,n}$ is $\operatorname{\mathrm{Proj}}$ of the semigroup algebra obtained from $C^{n}_{1r}$, the image of its integrable system is naturally seen as $P_{r,n}$.
@@
The finite set $\{f_{T^{\pm}}:T\text{ primitive semi-standard tableaux pair}\}$ is a SAGBI basis for the Cox ring of $\mathrm{K}^{n}_{r_{1}r_{2}}$, when $\gcd(r_{1},r_{2})=1.$
@@
In the next theorem, we show that $\mathrm{K}^{n}_{r_{1}r_{2}}$ has a toric degeneration to the toric variety associated to the intersection of (scaled) Gelfand–Cetlin polytopes $P_{r_{1},nr_{2}}$ and $P_{r_{2},nr_{1}}$. The lattice polytope $P_{r_{1},nr_{2}}$ lives in the toric lattice $M_{1}:=\operatorname{\mathrm{Mat}}(r_{1},nr_{2};\mathbb{Z})$, and the lattice polytope $P_{r_{2},nr_{1}}$ lives in the lattice $M_{2}:=\operatorname{\mathrm{Mat}}(r_{2},nr_{1};\mathbb{Z})$. We identify both of these lattices with $M=\operatorname{\mathrm{Mat}}(r_{2},r_{1};\mathbb{Z})^{\times n}$, by taking a tuple of matrices in $M$ and either stacking them horizontally (giving an element of $M_{2}$) or vertically and then transposing (giving an element of $M_{1}$), so that taking an intersection makes sense.
@@
We claim that
@@
In general, $P^{n}_{r_{1},r_{2}}$ is not a lattice polytope, though there is some smallest positive integer $k$ such that $kP^{n}_{r_{1},r_{2}}$ is a lattice polytope. This corresponds to the fact that the SAGBI basis is not generated in degree 1. It would be interesting to determine bounds on $k$.